# CHARME: A chain-based reinforcement learning approach for the minor embedding problem

### v1

@@ 4.2.2. The performance of CHARME
Figure 8. The frequency of finding the best solutions in term of qubit usage of three methods applied to three groups of the real world dataset categorized as (a) low density (b) medium density (c) high density.
The real dataset. Here, we assess the performance of CHARME in comparison to OCT-based and Minorminer in a real dataset. To evaluate these methods, we measure the frequency at which each approach achieves the best solutions across three testing groups: low density, medium density, and high density. Additionally, we analyze the running time of the three methods for various sizes of graphs.
Figure 8 illustrates the frequency of finding the best solutions in term of qubit usage of three methods for various testing groups including low density, medium density and high density. First, the CHARME method consistently outperforms the competitors, achieving the highest frequency of finding solutions with the fewest qubits in all three testing groups. Remarkably, the frequency of CHARME method is always above 50%, illustrating its superiority. Another interesting observation is in the performance of the OCT-based method which works well on the synthetic dataset, but struggles on the actual QUBO-based dataset. This phenomena is due to the sparse nature of logical graphs formed by actual QUBO formulation, making top-down approaches like Minorminer and CHARME more efficient choices.
Figure 9. Comparison between three methods in term of running time in the real world dataset with the number of nodes ranging from 6 to 165.
On the other hand, Figure 9 presents the running time analysis of the three methods. We observe a similar trend in the running time for the real-world dataset compared to the synthetic dataset. Specifically, the running time for Minorminer and CHARME method is comparable, while the OCT-based method exhibits significantly higher running times. That further confirms the efficiency of top down approaches in handling the real world dataset. In addition, the running time line of CHARME is smoother than lines of Minorminer and OCT-based which exhibit some fluctuations. This observation suggests that CHARME’s performance is more stable compared to its competitors.
@@ 5. Conclusion
In conclusion, our work introduces CHARME, a novel approach using Reinforcement Learning (RL) to tackle the minor embedding Problem in QA. Through experiments on synthetic and real-world instances, CHARME  demonstrates superior performance compared to existing methods, including fast embedding techniques like Minorminer and ATOM, as well as the OCT-based approach known for its high-quality solutions but slower runtime. Additionally, our proposed exploration strategy enhances the efficiency of CHARME’s training process. These results highlight the potential of CHARME  to address scalability challenges in QA, offering promising avenues for future research and application in quantum optimization algorithms.

### v2

@@ 4.2.3. The Performance of CHARME
Figure 10. The frequency of finding the best solutions and the corresponding average qubit usage (shown as numbers in parentheses)
of three methods applied to three groups of the real world dataset categorized as (a) low-density (b) medium-density (c) high-density.
The Real Dataset. Here, we assess the performance of CHARME in comparison to OCT-based and Minorminer on a real dataset. To evaluate these methods, we measure the frequency at which each approach achieves the best solutions across three testing groups: low-density, medium-density, and high-density. Additionally, we analyze the running time of the three methods for various sizes of graphs. We note that in this experiment, ATOM, ATOM (random), and ATOM (degree) do not achieve the best solution in any case. Thus, in order to avoid redundancy, we do not include these methods in this experiment.
Figure 10(c) illustrates the frequency of finding the best solutions in terms of qubit usage of three methods for various testing groups including low-density, medium-density, and high-density. First, the CHARME method consistently outperforms competitors, achieving the highest frequency of finding solutions with the fewest qubits in the three testing groups. Remarkably, the frequency of CHARME is always above 50%, illustrating its superiority. Another interesting observation is in the performance of the OCT-based method, which works well on the synthetic dataset, but struggles on the actual QUBO-based dataset. This phenomenon is due to the sparse nature of logical graphs formed by actual QUBO formulations, making bottom-up approaches like Minorminer and CHARME more efficient choices. In addition, the average qubit usage for the best solutions found by CHARME is higher than that found by Minorminer for low-density graphs. This suggests that CHARME performs better on low-density graphs with a large number of nodes, whereas Minorminer is more effective for graphs of a smaller size. However, this trend reverses for medium and high-density graphs, where CHARME excels on smaller graphs, and Minorminer proves to be more effective on larger ones.
Figure 11. Comparison between three methods in terms of running time in the real world dataset with the number of nodes ranging from 6 to 165.
On the other hand, Figure 11 presents the running time analysis of the three methods. We observe a similar trend in the running time for the real-world dataset compared to the synthetic dataset. Specifically, the running time for Minorminer and CHARME is comparable, while the OCT-based method exhibits significantly higher running times. That further confirms the efficiency of bottom-up approaches in handling the real-world dataset. In addition, the running time curve of CHARME is smoother than those of Minorminer and OCT-based, which exhibit fluctuations. This observation suggests that CHARME’s performance is more stable compared to its competitors.
@@ 5. Conclusion
In conclusion, our work introduces CHARME, a novel approach using Reinforcement Learning (RL) to tackle the minor embedding problem in Quantum Annealing (QA). Through experiments on synthetic and real-world datasets, CHARME demonstrates superior performance compared to existing methods, including fast embedding techniques such as Minorminer and ATOM, as well as the OCT-based approach known for its high-quality solutions but slower runtime. Additionally, our proposed exploration strategy enhances the efficiency of CHARME’s training process. These results highlight the potential of CHARME to address the scalability challenges in QA, offering promising avenues for future research and application in quantum optimization algorithms.